OpenStudy (anonymous):

There are 5 lollipops. Once a day, Joe licks on only one of them per day, and does this for 5 days straight. What is the probability that Joe did not lick on exactly 1 of the lollipops?

OpenStudy (anonymous):

that is way way wrong. nvm

OpenStudy (anonymous):

have to think more

OpenStudy (anonymous):

I got \[(1 \times \frac{4}{5} \times \frac{3}{5} \times \frac{2}{5})(\frac{4}{5} = \frac{96}{625})\]

OpenStudy (anonymous):

Sorry I put the parenthesis in the wrong place, but nonethless 96/625

OpenStudy (anonymous):

that looks much better than what i wrote

OpenStudy (anonymous):

Why is math so hard

OpenStudy (anonymous):

these problems are hard to wrap your mind around

OpenStudy (anonymous):

The second time he chooses to lick a lollipop, does he still have 5 lollipops to choose from or is it 4?

OpenStudy (anonymous):

still thinking lets put one lollipop aside and then see if we can figure the probability that he picks each one of the other 4

OpenStudy (anonymous):

Condition 1) 1st: 5 choices 2nd: 4 choices 3rd: 3 choices 4th: 2 choices 5th: 4 choices Therefore \[\frac{5 \times 4^2 \times 3 \times 2}{5^5}\]

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

Wow, are you joking? I was typing a long response and I clicked 'Back' by accident and it erased my entire reply

OpenStudy (anonymous):

annoying isn't it?

OpenStudy (anonymous):

Well, I guess I will just guess on this question

OpenStudy (anonymous):

i feel useless, but i am sure this is not that hard

OpenStudy (anonymous):

I'm the worst math major ever

OpenStudy (anonymous):

you should have seen me!

OpenStudy (anonymous):

i think your answer looks good though is this one of those annoying "sterling number" problems?

OpenStudy (anonymous):

*stirling